GGF4LIfeAlways:

Which of the following statements have the same result? Explain each step in solving each one. I. f(3) when f(x) = 2x + 2 II. f−1(4) when f(x) =3x minus 4, all over 5 III. y + 10 = 2y + 1

GGF4LIfeAlways:

Please help,it is very much needed at the moment.

GGF4LIfeAlways:

Anyone out there.........?I have been waiting for quite some time.

Vocaloid:

for I: f(3) means plug in x = 3 into the function. what does 2x + 2 equal when x = 3? II: it's asking for the value of the inverse function at x = 4. so you'll need to calculate the inverse of f(x). f(x) =(3x - 4)/5 swap the f(x) and the x to get x = (f(x) - 4)/5 and solve this to get f(x). once you do that, plug x = 4 into f(x). III: y + 10 = 2y + 1 solve this equation. you can subtract y from both sides, then subtract 1 from both sides, to solve for y. once you have all three values calculated, see which ones have the same answer.

Vocaloid:

correcting a small typo: f(x) =(3x - 4)/5 swap the f(x) and the x to get x = (3f(x) - 4)/5 needs to be a 3 in front of the f(x)

GGF4LIfeAlways:

Vocaloid ,I want to say thank you for your time on helping me on my question.Overall, I am very appreciative of you help.Is that as well all the help and or steps for each function,to be exact?

Vocaloid:

yes, these are the steps for the calculations for each function. is anything unclear?

GGF4LIfeAlways:

No,just making sure.Have a good rest of your day,Vocaloid.By the way,I am a big fan of your profile picture.